40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1652     Entry time: Sun Jun 26 21:44:06 2016
Author: awade 
Type: DailyProgress 
Category: Laser 
Subject: South laser mode hopping 

In a previous post (PSL_Lab/1641) I identified a possible mode hop not far from our typical operating point by tuning the laser diode current down two clicks from the typical Adj# = 0 point. It is likely that the slightly lower temperature/operating point of the laser diode put it closer to the edge.

The experiment was intermittently dropping lock last time it was in operation and it is likely that mode hopping in the south laser was a culprit.

I tried reproducing this by putting a 2 Vpp ramp on the slow temperature control input of the laser controller.  I was able to reproduce some clear mode hopping.  This behavior didn't really set in until the laser had had some time to warm up.  It seemed less prevalent at the default factory temperature set point of 48 C, but our operating point was closer to 34.4965 C (this is to match the north laser). At one point the mode hopping was close to the V_slowcontrols = 0 but after turning the laser off and on again I couldn't find this.

The added difficulty is that both lasers must be kept to within the beat note detection bandwidth and so we must find a sweet spot of no mode hopping between the two lasers.  Tomorrow, after some warm up, I will try some temperatures around 34.5 C to see if there is a sweet spot. Need to confirm with Aidan what the intended slow controls range is.  From the manual we know we should get about 3.1GHz/˚C with a mode hop at intervals of >10 GHz.  We don't need this much range but need to make sure the laser has the stable region appropriately centered.

Pictures all at the 34.4965 C setpoint after turning the laser off for a period and then back on

Attachment 1: What scan looks like shortly after turning on laser (0.01 Hz ramp + PD signal picked off from the laser path)

Attachment 2: Discontinous jump induced by temperature ramp (0.01 Hz ramp + PD signal picked off from the laser path)

Attachment 3: Some hoping, was much more distinctive earlier in the evening (0.01 Hz ramp + PD signal picked off from the laser path)

Sorry these look like they were taken with a potato, I didn't have a USB drive and phone was all I had.

 

Attachment 1: 2016-06-26_21.16.42.jpg  793 kB  | Hide | Hide all
2016-06-26_21.16.42.jpg
Attachment 2: 2016-06-26_21.16.32.jpg  794 kB  | Hide | Hide all
2016-06-26_21.16.32.jpg
Attachment 3: 2016-06-26_21.15.17.jpg  804 kB  | Hide | Hide all
2016-06-26_21.15.17.jpg
ELOG V3.1.3-